which element's 4+ ion would have the following electron configuration?

[Ar]3d^10
Answer #1

[Ar]3d10

Ar = 18

3d having 10

total electrons = 18 + 10 = 28  

+4 ion means if we add 4 electrons we get neutral atom

28 + 4 = 32 electrons

so neutral atom contain 32 electrons

for neutral atom number of electrons = atomic number = 32

atomic number 32 = Germanium = Ge

answer = Germanium = Ge